1 1 2 -1] TO 1 01 0 0] 7. The product PFA = 2 5 3 where P = 0 0 1 and F = 0 1-1 9 4. 11 00 .-1 0 11 matrix A. 1 Find the
Answer #1

& The product २ TI -) where p = PFA 3 5 2 J -1 9 0 Find Matrix A. 0 and F = 0 0 Sol Let Ba २ 5 3 2 - 1 9 Thus = B Puy by pt Bp or 0 0 pl 11 0 - FE 0 Fs - 0 1 Thus A = F-prB 2 0 3 A- 5 9 11:34 ima o 4 1 3X3 2 - - A = o 2 2 5 9 3 4 4 -1 A- 9 २ 14 -) 7